OpenStudy (christos):

Find the exact value of -sec10*/csc80*

OpenStudy (anonymous):

\[\frac{ -(\sec(10°)) }{ (\csc(80°))}\] \[-1\]

OpenStudy (anonymous):

-

OpenStudy (anonymous):

@Jezebel493 @Christos ?

OpenStudy (christos):

But how did you go to that solution? @some_someone

OpenStudy (anonymous):

?

OpenStudy (anonymous):

how did you go to that solution

OpenStudy (anonymous):

some_someone?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

out

OpenStudy (anonymous):

Christos

OpenStudy (christos):

please help me anyone

hartnn (hartnn):

again ? using the identity \(\csc x=\sec(90-x)\) if we put x =80, we get, \(\csc 80=\sec(90-80)=\sec10\) hence, -sec10/csc 80 = -sec 10/ sec 10 = -1

OpenStudy (anonymous):

Good, that is what I was going to reply right now. I could not earlier cause i was not online but yeah @hartnn Great Job :P